The rate of diffusion of a same of ozonised oxygen is 0.98 times more than that of pure oxygen. Find the percentage (by volume) of ozone in the ozonised sample.

Grade:11

1 Answers

Vikas TU
6 years ago
We realize that dissemination rate is conversely relative to the square foundation of the molar masses. 
What's more, ozonized gas diffuses slower than O2 then it has higher molar mass since ozone is O3 
The molar weight of ozonized gas =(32g/mole)/(0.982) 
=33.32g/mol 
The blend contains O2 with molar mass 32g/mole and Ozone with molar mass of 48g/mole 
On the off chance that we let X be the mole portion of O2 and 1-X be the mole division of Ozone. 
At that point, 
X*32g/mole +(1-X)*48g/mole =33.32g/mole 
X=14.68/16 
Rate of ozone; 
=(1-X)*100% 
=(1-14.68/16)*100% 
Ozone =8%.